# Representative Action Selection for Large Action-Space Meta-Bandits

### v1

@@ Appendix H Proof of Theorem 8
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 14. The second inequality follows from Equation (19). The third inequality uses the Pólya-Szegő inequality [17], corresponding to a constant $c>0$:
@@
Appendix I Example of orthogonal basis
In general, even when the distribution of bandit instances $\theta$ is well-behaved, e.g., the standard Gaussian distribution in $\mathbb{R}^{n}$, the importance measure $q$ is usually non-trivial. In this section, we consider an example of the action space being orthogonal basis.
Let the set $\mathcal{A}_{full}=\{e_{i}:i=1,\dots,n\}$ consists of the $n$ unit vectors aligned with the coordinate axes in $\mathbb{R}^{n}$. Hence, $\operatorname{diam}(\mathcal{A}_{full})=\sqrt{2}$. In this case, $\max_{a\in\mathcal{A}_{full}}\langle a,\theta\rangle$ is equivalent to the maximum among the $n$ entries of $\theta$, where by definition, each entry follows i.i.d. from the standard normal distribution $\mathcal{N}(0,1)$, such that each entry has the same probability of having the maximum value. Let each cluster contain only one unit vector, thus the importance measure of clusters $q$ is a uniform distribution.
Expected maximum in $\mathcal{A}_{full}$: Let $X_{i}$, $i=1,\dots,n$ be i.i.d. samples from $\mathcal{N}(0,1)$, then
@@ Appendix I Example of orthogonal basis
$\mathbb{E}\max_{a\in\mathcal{A}_{full}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
@@
$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{full}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
@@
Upper bound of expected regret of Algorithm 1: Since the importance measure of clusters $q$ is a uniform distribution, we have
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\left(1-\frac{1}{n}\right)^{K}\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right]\leq\left(1-\frac{1}{n}\right)^{K}\sqrt{2\log n},$
where the left inequality uses Corollary 16 and that if each cluster contains one action, $\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]=0$ for each cluster and $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)=0$.
Corollary 16.
Let $\mathcal{A}$ be output of Algorithm 1. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, where the $q$ measure is uniform over all clusters. Let $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Then, for the same constant $C>0$ in Theorem 5, it holds that
$\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}+\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right].$
Under Assumption 4,
$\mathbb{E}_{\theta,\mathcal{A}}[\operatorname*{Regret}]\geq\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\left(\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}\right).$
Proof.
Consider the variables $Y_{\ell}$ defined in the proof of Theorem 6,
$\mathbb{E}[\operatorname*{Regret}]\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\sum_{r\in\mathcal{R}}q(r)\cdot\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\Big{|}a^{*}(\theta)\in r\right]$
$=$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{full}}\mu_{a}\right],$
where the first inequality comes from the proof of Theorem 6, the equality uses $q(r)=1/|\mathcal{R}|$ for all $r\in\mathcal{R}$. Then, using Equation (11) in the proof of Theorem 5, we derive the upper bound.
Consider the variables $Y_{\ell}$ defined in the proof of Theorem 8,
$\mathbb{E}[\operatorname*{Regret}]\geq$ $\sum_{\ell\leq m}q(r_{\ell})\cdot\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]$
$=$ $\left(1-\frac{1}{|\mathcal{R}|}\right)^{K}\cdot\sum_{\ell\leq m}q(r_{\ell})\cdot\mathbb{E}\left[Y_{\ell}\Big{|}a^{*}(\theta)\in r_{\ell}\right]\geq\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right],$
where the first inequality comes from the proof of Theorem 8, the equality uses $q(r)=1/|\mathcal{R}|$ for all $r\in\mathcal{R}$. The last inequality uses the definition of $q$ measure and $Y_{\ell}\geq\min_{\ell\leq m}Y_{\ell}$. Then, using Equation (12) in the proof of Theorem 5, we derive the lower bound. ∎
Appendix J Comparisons with UCB & TS
To understand why Algorithm 1 outperforms UCB and TS, we still consider the fixed action space of 15 grid points over the interval $[0,2]$ and indexed by $0,1,\ldots,14$. Outcome functions are generated using an RBF kernel with a fixed length-scale $l=1$. The middle upper subplot of Figure 4 displays ten example outcome functions sampled from this setup. In each trial, Algorithms 1, 2 (UCB), and 3 (TS) are run to select $K=5$ distinct representative actions. UCB and TS are each executed for 3000 rounds (excluding initialization), with each round corresponding to a bandit instance. The number of rounds is chosen to approximately match the number of possible super arms, given by $\binom{15}{5}=3003$.
The upper part of Figure 3 visualizes the representative actions selected by each algorithm in three individual trials. Each row corresponds to one trial. The lower part shows the frequency with which each of the 15 actions is selected across 50 independent trials. The frequencies for actions indexed 1 through 13 are relatively uniform across algorithms, but a notable distinction is that Algorithm 1 (green) exhibits a strong preference toward the boundary actions—those indexed 0 and 14.
To understand this preference of Algorithm 1, recall that its output frequencies directly reflect the underlying measure $q$, since the algorithm samples actions according to $q$ by design. In Figure 4, we vary the RBF kernel length-scale across values 0.1, 1, and 4. For each setting, we sample ten example outcome functions (shown in the upper part) and run Algorithm 1 for 50 trials, recording the frequency of each selected action as an approximation of the $q$-measure (shown in the lower part).
In Figure 4, when the length-scale is 4, the outcome functions are very smooth—almost linear—so the maximum is typically attained at the boundary actions (indices 0 and 14), causing the $q$-measure to concentrate heavily on them. In contrast, with a length-scale of 0.1, the functions are very rough, with outcomes nearly independent across actions, resulting in an approximately uniform $q$-measure over the action space. Further, since the outcome functions are generated using a stationary RBF kernel, the corresponding feature representations $\Phi(a)$, for $a\in\mathcal{A}_{full}$, are uniformly distributed. When action outcomes are nearly independent (e.g., with a very small length-scale), the feature vectors $\Phi(a)$ form an orthogonal basis, with each pair $(\Phi(a),\Phi(a^{\prime}))$ equidistant. As discussed in Supplementary I, this leads to a uniform $q$-measure when each cluster contains a single action. Indeed, the lower-left part of Figure 4 confirms this: for length-scale $l=0.1$, the action selection frequencies are nearly uniform. As the length-scale increases, action outcomes become more correlated, and the pairwise distances between $\Phi(a)$ vectors shrink. The boundary actions (indices 0 and 14), remain the farthest apart and thus are more likely to be optimal in a given bandit instance.
This explains why Algorithm 1 performs better: it closely aligned with the $q$-measure, whereas UCB and TS select actions more uniformly. Further, in any given round $\theta$ of TS/UCB, the reward $\max_{a\in\mathcal{A}}\mu_{a}$ for optimal and suboptimal super arms can be indistinguishable if both include the optimal base action $a^{*}(\theta)$. This ambiguity slows down the convergence of the learning algorithms because they cannot effectively differentiate between good and bad super arms—a problem exacerbated by the combinatorial number of possible super arms.
Figure 3: Comparison of representative actions selected by three algorithms in a synthetic setting. Outcome functions are generated using an RBF kernel (length-scale $l=1$) over 15 grid points in $[0,2]$. In each trial, $K=5$ distinct actions are selected per algorithm (Algorithm 1 in green, TS in yellow, and UCB in red); UCB and TS are run for 3000 rounds. Upper: Representative actions selected in three individual trials; each row shows one trial. Lower: Action selection frequencies of Algorithm 1, TS, and UCB over 50 trials. Algorithm 1 consistently selects boundary actions (indices 0 and 14), while UCB and TS select more uniformly.
Figure 4: Effect of kernel length-scale on importance measure $q$. Upper: Example outcome functions generated using RBF kernels with varying length-scales $l\in\{0.1,1,4\}$, over 15 grid points in $[0,2]$, with actions indexed from 0 to 14. Lower: Action selection frequencies of Algorithm 1 over 50 trials for each length-scale setting. These frequencies approximate the importance measure $q$. As the length-scale increases, the outcome functions become smoother and more linear, causing $q$-measure to concentrate at the boundary actions (indices 0 and 14).
Algorithm 2 Upper Confidence Bound with Gaussian Rewards
1:Input: Base actions $\mathcal{A}_{full}$, subset cardinality $K$, kernel matrix $\mathbf{K}$, time horizon $T=3000$
2:Initialize: Super arms $\{\mathcal{A}_{i}\}_{i\leq M}$ with $M=\binom{|\mathcal{A}_{full}|}{K}$, reward variance $\sigma^{2}=1$
3:Initialize: $N_{i}\leftarrow 0$, $\hat{\mu}_{i}\leftarrow 0$ for all $i=1,\dots,K$
4:for $t=1$ to $T+M$ do $\triangleright$ Each round corresponds to a bandit
5:     Sample an outcome function $f\sim\mathcal{N}(0,\mathbf{K})$
6:     if $t\leq M$ then
7:         Select arm $i_{t}=t$
8:     else
9:         Select arm $i_{t}=\arg\max_{i}\left(\hat{\mu}_{i}+\sqrt{\frac{2\log t}{N_{i}+\epsilon}}\right)$
10:     end if
11:     Pull arm $i_{t}$, observe reward $r_{t}=\max_{a\in\mathcal{A}_{i_{t}}}f(a)$
12:     Update:
$N_{i_{t}}\leftarrow N_{i_{t}}+1$
$\hat{\mu}_{i_{t}}\leftarrow\hat{\mu}_{i_{t}}+\frac{r_{t}-\hat{\mu}_{i_{t}}}{N_{i_{t}}}$
13:end for
Algorithm 3 Thompson Sampling with Gaussian Priors
1:Input: Base actions $\mathcal{A}_{full}$, subset cardinality $K$, kernel matrix $\mathbf{K}$, time horizon $T=3000$
2:Initialize: Super arms $\{\mathcal{A}_{i}\}_{i\leq M}$ with $M=\binom{|\mathcal{A}_{full}|}{K}$, reward variance $\sigma^{2}=1$
3:Initialize: prior mean $\mu_{i}\leftarrow 0$, prior precision $\tau_{i}\leftarrow\epsilon$ for all $i=1,\dots,K$
4:for $t=1$ to $T$ do $\triangleright$ Each round corresponds to a bandit
5:     Sample an outcome function $f\sim\mathcal{N}(0,\mathbf{K})$
6:     for $i=1$ to $M$ do
7:         Sample $\tilde{\mu}_{i}\sim\mathcal{N}(\mu_{i},\sigma^{2}/\tau_{i})$
8:     end for
9:     Select arm $i_{t}=\arg\max_{i}\tilde{\mu}_{i}$
10:     Pull arm $i_{t}$, observe reward $r_{t}=\max_{a\in\mathcal{A}_{i_{t}}}f(a)$
11:     Update:
$\tau_{i_{t}}\leftarrow\tau_{i_{t}}+1$
$\mu_{i_{t}}\leftarrow\mu_{i_{t}}+\frac{r_{t}-\mu_{i_{t}}}{\tau_{i_{t}}}$
12:end for
@@
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [28]:

### v2

@@ Appendix F Proof of Theorem 11
where the equality uses the tower rule. The first inequality holds because the regret is lower bounded by zero when $r\cap\mathcal{A}\neq\emptyset$ and $a^{*}(\theta)\in r$, and Lemma 17. The second inequality follows from Equation (19). The third inequality uses the Pólya-Szegő inequality [Dragomir, 2004], corresponding to a constant $c>0$:
@@
Appendix G Proof of Theorem 12
Definition 2.
The random process $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a mean-zero sub-Gaussian process if the process $\mathbb{E}\mu_{a}=0$ and has the increment condition:
$\forall u>0,\Pr[|\mu_{a}-\mu_{a^{\prime}}|\geq u]\leq 2\exp\left(-\frac{u^{2}}{2\|a-a^{\prime}\|^{2}_{2}}\right).$
Let $\gamma_{2}(\mathcal{S})$ be the Talagrand’s chaining functional [Talagrand, 2014] for a set $\mathcal{S}\in\mathbb{R}^{n}$, $n\in\mathbb{N}\cup\{+\infty\}$ and Euclidean norm. It is well-known that under the assumption that $\{\mu_{a}\}_{a\in\mathcal{S}}$ is a Gaussian process, this gives the tightest bound, for a universal constant $L$:
$\frac{1}{L}\gamma_{2}(\mathcal{S})\leq\mathbb{E}\left[\max_{a\in\mathcal{S}}\mu_{a}\right]\leq L\gamma_{2}(\mathcal{S}).$
Within the proof of the upper bound for general sub-Gaussian process, it also derives the deviation bound for the term $\max_{a\in\mathcal{S}}\mu_{a}$:
Lemma 19 (Theorem 2.2.22 of Talagrand [2014]).
Let $\{\mu_{a}\}_{a\in\mathcal{S}}$ be a mean-zero sub-Gaussian process, then there exists a constant $c>0$:
$\Pr\left[\sup_{a,a^{\prime}\in\mathcal{S}}|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]\leq 2\exp\left(-\frac{cu^{2}}{\gamma^{2}_{2}(\mathcal{S})}\right).$
Lemma 20.
Let $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, and an arbitrary reference subset $\mathcal{A}:=\{a_{\ell}\}_{\ell\leq m}$.
Fix $\ell$, define a non-negative random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Then, for some constant $C>0$,
$\mathbb{E}_{\theta}\left[\max_{\ell\leq m}Y_{\ell}\right]\leq C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell}).$
Proof.
Let $\epsilon:=\max_{\ell\leq m}\gamma_{2}(r_{\ell})$. For any $\ell\leq m$, we have
$\Pr\left[|Y_{\ell}|\geq u\right]=$ $\Pr\left[\left|\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right|\geq u\right]$
$\leq$ $\Pr\left[\sup_{a,a^{\prime}\in r_{\ell}}|\mu_{a}-\mu_{a^{\prime}}|\geq u\right]$
$\leq$ $2\exp\left(-\frac{cu^{2}}{\gamma^{2}_{2}(r_{\ell})}\right)$
$\leq$ $2\exp\left(-\frac{cu^{2}}{\epsilon^{2}}\right),$
where the equality uses definition of $Y_{\ell}$. The first inequality uses
$\left|\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}\right|\leq\sup_{a\in r_{\ell}}|\mu_{a}-\mu_{a_{\ell}}|\leq\sup_{a,a^{\prime}\in r_{\ell}}|\mu_{a}-\mu_{a^{\prime}}|.$
The second inequality uses the fact that $\{\mu_{a}\}_{a\in r_{\ell}}$ is a mean-zero sub-Gaussian process and applies Lemma 19. The last inequality uses the definition of $\epsilon$, and $c>0$ is a constant.
By union bound, we have
$\Pr\left[\max_{\ell\leq m}|Y_{\ell}|\geq u\right]\leq 2m\exp\left(-\frac{cu^{2}}{\epsilon^{2}}\right).$
Further, using Lemma 14, we have for some absolute constant $C>0$:
$\mathbb{E}\left[\max_{\ell\leq m}|Y_{\ell}|\right]\leq C\epsilon\sqrt{\log m}.$
Statement of Theorem 12:
Let $\mathcal{A}$ be output of Algorithm 1. Let $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ be a mean-zero sub-Gaussian process. Consider a partition :=$\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space. Then, for the same constant $C>0$,
$\begin{split}\mathbb{E}_{\theta,\mathcal{A}}[\textnormal{Regret}]\leq&C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})\\ &+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}.\end{split}$
Proof.
If $r_{\ell}\cap\mathcal{A}\neq\emptyset$, define $a_{\ell}\in r_{\ell}\cap\mathcal{A}$. If $r_{\ell}\cap\mathcal{A}\!=\!\emptyset$, choose an arbitrary point $a_{\ell}\in r_{\ell}$ as the representative. The set $\mathcal{A}^{\prime}:=\{a_{\ell}\}_{\ell\leq m}$ forms a reference subset of Definition 1.
For each $\ell\leq m$, define a random variable
$Y_{\ell}:=\sup_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}.$
Following the same reasoning used in the proof of Theorem 8:
@@ Appendix G Proof of Theorem 12
$\mathbb{E}[\operatorname*{Regret}]\leq$ $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2}$
$\leq$ $C\sqrt{\log m}\cdot\max_{\ell\leq m}\gamma_{2}(r_{\ell})+\left(\mathbb{E}_{q}\left[(1-q(r))^{2K}\right]\cdot\mathbb{E}_{\theta}\left[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu^{2}_{a}\right]\right)^{1/2},$
The first inequality follows from Equation (14) in the proof of Theorem 8. The last inequality uses Lemma 20. ∎
Appendix H The effect of clustering structure on regret
Appendix I Varying-dependence actions with RBF/Gibbs kernels
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel [Williams and Rasmussen, 2006].
[…]
where $l$ is a length-scale parameter and $l(a):=0.1+0.9\cdot\exp(-\|a\|^{2})$ is a location-dependent length-scale function. Both of them control the dependence between actions. But, unlike the stationary RBF kernels, the Gibbs kernel allows the correlation to depend not only on the distance between actions, but also on their locations. When $l(a)=l$ is a constant, the Gibbs kernel reduces to the RBF kernel.
Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, depending on the choice of kernel. We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$. Under either kernels defined in Equation (20), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{\operatorname*{full}}$. In this way, we sample functions from a RKHS function class; See [Kanagawa et al., 2018, Theorem 4.12].
Figure 3: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (20). Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in [0,2][0,2][ 0 , 2 ] (blue curves, right y-axis). The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points.
To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval [0,2][0,2][ 0 , 2 ], using Gibbs kernel in Equation (20) to sample outcome functions. To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid. Figure 3 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region. We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected. The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart. Another interesting aspect of this subplot is the two high bars at the edges. Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$. For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes. The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
Appendix J Independent and identically distributed actions
As an extreme case, suppose that $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ is a set of i.i.d. random variables. This corresponds to the canonical process in which the action space is given by the orthonormal basis of $\mathbb{R}^{n}$, where $n=|\mathcal{A}_{\operatorname*{full}}|$. To see this, we associate each action $a$ with a unit vector $e_{a}$, which has a value of 111 at the $a$th coordinate and 0 elsewhere, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$. With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ consists of mutually independent random variables.
Let the set $\mathcal{A}_{\operatorname*{full}}=\{e_{i}:i=1,\dots,n\}$ denote the $n$ unit vectors aligned with the coordinate axes in $\mathbb{R}^{n}$. Hence, $\operatorname{diam}(\mathcal{A}_{\operatorname*{full}})=\sqrt{2}$. In this case, $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$ is equivalent to the maximum among the $n$ entries of $\theta$, where each entry is i.i.d. from the standard normal distribution $\mathcal{N}(0,1)$. By symmetry, each coordinate has the same probability of attaining the maximum value. If each cluster contains only one unit vector, then the importance measure $q$ over clusters is the uniform distribution.
Expected maximum in $\mathcal{A}_{\operatorname*{full}}$: Let $X_{i}$, $i=1,\dots,n$ be i.i.d. samples from $\mathcal{N}(0,1)$, then
$\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}=\mathbb{E}_{\theta}\max_{i=1,\dots,n}\theta_{i}=\mathbb{E}\max_{i=1,\dots,n}X_{i},$
@@ Appendix I Varying-dependence actions with RBF/Gibbs kernels
$\frac{\sqrt{\log n}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\sqrt{2\log n},\quad\frac{\sqrt{\log|\mathcal{A}|}}{\sqrt{\pi\log 2}}\leq\;\mathbb{E}\max_{a\in\mathcal{A}}\mu_{a}\leq\sqrt{2\log|\mathcal{A}|}.$
@@
Let $X_{1},\dots,X_{N}$ be i.i.d. $\mathcal{N}(0,\sigma^{2})$ random variables, then according to [Kamath, 2015]: